A square has a perimeter of 80 meters. what is its area? Please answer with a representation and checking.
Answer by mananth(16946)

Square - all sides are equal
perimeter = 80
side length = 80/4 = 20m
Area = side ^2
area = 20*20 = 400 sq.m